### v2

# OBLR-PO: A Theoretical Framework for Stable Reinforcement Learning

Abstract
Existing reinforcement learning (RL)-based post-training methods for large language models have advanced rapidly, yet their design has largely been guided by heuristics rather than systematic theoretical principles. This gap limits our understanding of the properties of the gradient estimators and the associated optimization algorithms, thereby constraining opportunities to improve training stability and overall performance. In this work, we provide a unified theoretical framework that characterizes the statistical properties of commonly used policy-gradient estimators under mild assumptions. Our analysis establishes unbiasedness, derives exact variance expressions, and yields an optimization-loss upper bound that enables principled reasoning about learning dynamics. Building on these results, we prove convergence guarantees and derive an adaptive learning-rate schedule governed by the signal-to-noise ratio (SNR) of gradients. We further show that the variance-optimal baseline is a gradient-weighted estimator, offering a new principle for variance reduction and naturally enhancing stability beyond existing methods. These insights motivate Optimal Baseline and Learning-Rate Policy Optimization (OBLR-PO), an algorithm that jointly adapts learning rates and baselines in a theoretically grounded manner. Experiments on Qwen3-4B-Base and Qwen3-8B-Base demonstrate consistent gains over existing policy optimization methods, validating that our theoretical contributions translate into practical improvements in large-scale post-training. * Equal Contribution.
{zyzheng,alexpku,jiayi_sheng}@berkeley.edu
1 Introduction
Reinforcement learning (RL) has become a central paradigm for training large-scale models to exhibit complex reasoning and decision-making abilities [7, 33, 24, 2, 11, 32]. Policy optimization algorithms such as Proximal Policy Optimization (PPO) [29], Group Relative Policy Optimization (GRPO) [11, 30], and REINFORCE with leave-one-out baselines (RLOO) [1] explore different designs of reward modeling and advantage estimation to guide learning. Despite these advances, training stability remains a fundamental bottleneck, particularly in the post-training stage of large language models [6, 27, 41, 34].
Training stability is influenced not only by algorithmic design but also by choices such as learning rate schedules and baseline functions. In large-scale pre-training, adaptive or decayed learning rate schedules are routinely used to ensure stable optimization and steady convergence [3, 4, 10, 19, 9, 18]. It is therefore natural to ask whether similar benefits could be realized in reinforcement learning, where instability remains a persistent challenge [37]. Baseline design has likewise been explored as a practical tool for variance reduction, with numerous heuristics proposed across different policy optimization methods [43, 1, 35]. However, most of these strategies are empirical in nature, and the field still lacks systematic analysis that clarifies what constitutes an effective baseline and how it influences learning dynamics.
These limitations highlight the need for deeper theoretical guidance. While empirical techniques for stability are widely adopted, systematic theory for post-training policy optimization is still scarce [25, 40, 5, 39, 20]. Under mild and interpretable assumptions, rigorous analysis can illuminate the statistical properties of gradient estimators, reveal how baseline choices interact with learning dynamics, and clarify when adaptive learning rates can provably improve stability. Such theoretical insights not only fill a long-standing gap in understanding but also directly motivate algorithmic designs that bridge principled analysis with practical effectiveness.
To validate our theoretical findings, we implement the proposed algorithm on Qwen3-4B and Qwen3-8B models. Across multiple benchmarks, our method consistently outperforms existing policy optimization baselines, demonstrating both improved stability and stronger performance. These results confirm that the theoretically motivated learning rate schedule and baseline design translate into tangible gains in practical large-scale post-training.
Our main contributions are as follows:
  • We present a unified theoretical framework for policy optimization, and under mild assumptions derive unbiasedness, variance expressions, and an upper bound on the optimization loss (Section 3, Section 4.1, Section 4.2).
  • We optimize this upper bound to obtain the optimal learning rate schedule, governed by the gradient signal-to-noise ratio (Section 4.3).
  • We characterize the optimal baseline design, showing that a gradient-weighted form achieves principled variance reduction (Section 4.4).
  • We propose the Optimal Baseline and Learning-Rate Policy Optimization (OBLR-PO) algorithm and empirically validate its stability and performance improvements on Qwen3-4B-Base and Qwen3-8B-Base over existing policy optimization methods (Section 5, Section 6).
2 Related Work
Theoretical Foundations of Policy Optimization
Recent studies have analyzed the training dynamics of policy optimization, especially focusing on the loss function upper bound and the resulting convergence guarantees [20, 25, 5, 39]. For example, the impact of a single-step update on the loss, along with the identification of the optimal update vector, has been analyzed [20]. Similarly, stochastic no-regret oracle frameworks have been employed to provide theoretical guarantees, leading to regret upper bounds and formal connections to online learning [5]. In parallel, classical gradient descent optimization theory has been applied to establish convergence rates under smoothness and convexity assumptions [25, 39]. Related efforts also bridge theory and practice by providing convergence results for GRPO and related algorithms, thereby supporting their empirical success [25]. Moreover, under smoothness conditions, it has been shown that the loss function admits a guaranteed decreasing rate [39]. Building on this analysis, our work establishes tighter upper bounds, provides stronger and more general convergence guarantees, and introduces principled strategies for learning rate schedules and baseline design.
Algorithmic Variants of Policy Optimization
Policy optimization, originating from reinforcement learning, is central to shaping the reasoning capabilities of large language models and has therefore become the foundation of reinforcement learning from human feedback (RLHF) [7, 33, 24, 2, 11]. Within this framework, supervised fine-tuning (SFT) [24, 36] provides initial alignment through imitation on instruction data, PPO [29] extends this paradigm with critic-based reinforcement learning on preference-model rewards, and Direct Preference Optimization (DPO) [26] further simplifies the objective by introducing a contrastive loss that directly matches human preferences between responses. To mitigate the complexity of critic-based methods, GRPO [11, 30] replaces the learned value function with a group-based baseline defined as the average reward within each candidate set, while alternative designs such as ReMax [21], RLOO [1], and Reinforce++ [16] adopt maximum-reward, leave-one-out, or variance-reduced baselines, respectively. Building on these developments, this work presents a general formulation that unifies the above algorithms under a common framework, enabling systematic theoretical analysis and leading to the identification of a principled optimal baseline.
3 Problem Setup
3.1 Objective Function
In this section, we formally define our problem setup. Our target is to learn an optimal policy $\pi_{\theta}$ that maximizes the expected reward, which serves as a measure of accuracy or performance on the given task. Formally, we aim to solve:
$\max_{\theta}\ \underbrace{\mathbb{E}_{q\sim D,o\sim\pi_{\theta}(\cdot|q)}\left[F(q,o)\right]}_{J(\theta)}.$ (1)
In online optimization algorithms, data is collected using an old policy $\pi_{\theta_{\text{old}}}$, and importance sampling is employed to correct for the discrepancy between the old policy and the target policy $\pi_{\theta}$ by scaling the advantage estimates with the importance sampling ratio $\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}$:
$\max_{\theta}\ \mathbb{E}_{q\sim D,o\sim\pi_{\theta_{\operatorname{old}}}(\cdot|q)}\left[\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\operatorname{old}}}(o|q)}F(q,o)\right].$ (2)
Here $D$ denotes the distribution over queries $q$, $\pi_{\theta}$ is a behavior policy which generates outputs $o$ conditioned on $q$. $F(q,o)$ is an approximate reward which takes the query $q$ and output $o$ and returns a scalar value representing the estimated quality of the output.
In our setup, the reward $F(q,o)$ is assumed to be available, either from accuracy supervision or a pre-trained reward model.
3.2 Related RL Algorithms
A variety of reinforcement learning algorithms have been developed to improve the reasoning ability of large language models. In this section, we present several representative methods, each formulated through a distinct surrogate objective $J_{\operatorname{PO}}(\theta)$.
PPO
The Proximal Policy Optimization (PPO) [29] objective is formulated as
$J_{\operatorname{PPO}}(\theta)=$ $\mathbb{E}_{q\sim D,\,o\sim\pi_{\theta_{\operatorname{old}}}(\cdot|q)}$
$\left[\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\operatorname{old}}}(o|q)}\,\hat{A}^{\mathrm{PPO}}(q,o)\right],$ (3)
where the advantage estimator $\hat{A}^{\operatorname{PPO}}$ is computed using Generalized Advantage Estimation (GAE) [28], a widely adopted variance-reduction technique that stabilizes policy-gradient training.
GRPO
Group Relative Policy Optimization (GRPO) [30] optimizes policies by comparing rewards among a group of sampled outputs, without relying on an explicit value or reward model. Given a query $q$, we draw $G$ outputs $\{o_{i}\}_{i=1}^{G}$ from the old policy $\pi_{\theta_{\operatorname{old}}}$ with associated rewards $r_{i}=F(q,o_{i})$. The objective is
$J_{\operatorname{GRPO}}(\theta)=$ $\mathbb{E}_{q\sim D,\,\{o_{i}\}\sim\pi_{\theta_{\operatorname{old}}}}$
$\left[\frac{1}{G}\sum_{i=1}^{G}\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\operatorname{old}}}(o_{i}|q)}\,\hat{A}^{\mathrm{GRPO}}(q,o_{i})\right],$ (4)
where the group-relative advantage is normalized as
$\hat{A}^{\mathrm{GRPO}}(q,o_{i})=\frac{r_{i}-\frac{1}{G}\sum_{j=1}^{G}r_{j}}{\sqrt{\frac{1}{G}\sum_{j=1}^{G}\left(r_{j}-\frac{1}{G}\sum_{k=1}^{G}r_{k}\right)^{2}}}.$ (5)
ReMax
The ReMax [21] method draws inspiration from the REINFORCE with Baseline approach, where we modify the gradient estimation by incorporating a subtractive baseline value. The objective is:
$J_{\operatorname{ReMax}}(\theta)=\mathbb{E}_{q_{i}\sim D,\,o^{i}_{1:T}\sim\pi_{\theta}(\cdot|q_{i})}$
$\left[\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\frac{\pi_{\theta}(o_{t}^{i}|q,o_{1:t-1}^{i})}{\pi_{\theta_{\operatorname{old}}}(o_{t}^{i}|q,o_{1:t-1}^{i})}\hat{A}^{\mathrm{ReMax}}(q_{i},o^{i}_{1:T})\right],$ (6)
where the action $o^{i}_{t}\sim\pi_{\theta}(\cdot|q_{i},o^{i}_{1:t-1})$, and $b_{\theta}(q_{i})$ is the baseline value. The choice for the baseline is:
$b_{\theta}(q_{i})=r(q_{i},\bar{o}^{i}_{1:T}),\quad\bar{o}^{i}_{t}\in\arg\max\pi_{\theta}(\cdot|q_{i},\bar{o}^{i}_{1:t-1}),$
This baseline value is obtained by greedily sampling the response and calculating the associated reward value.
The advantage function is defined as:
$\hat{A}^{\mathrm{ReMax}}(q_{i},o^{i}_{1:T})=r(q_{i},o^{i}_{1:T})-b_{\theta}(q_{i}).$
RLOO
REINFORCE Leave-One-Out (RLOO) [1, 17] extends the REINFORCE estimator to the multi-sample setting by employing a leave-one-out baseline. Given a query $q$, we draw $G$ outputs $\{o_{i}\}_{i=1}^{G}$ from the old policy $\pi_{\theta_{\operatorname{old}}}$ with rewards $r_{i}=F(q,o_{i})$. The objective is
$J_{\operatorname{RLOO}}(\theta)=$ $\mathbb{E}_{q\sim D,\,\{o_{i}\}\sim\pi_{\theta_{\operatorname{old}}}}$
$\left[\frac{1}{G}\sum_{i=1}^{G}\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\operatorname{old}}}(o_{i}|q)}\,\hat{A}^{\mathrm{RLOO}}(q,o_{i})\right],$ (7)
where the leave-one-out advantage is
$\hat{A}^{\mathrm{RLOO}}(q,o_{i})=r_{i}-\frac{1}{G-1}\sum_{j\neq i}r_{j}.$ (8)
General Form
The surrogate objective $J_{\operatorname{PO}}(\theta)$ can be expressed as
$J_{\operatorname{PO}}(\theta)=\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\operatorname{old}}}(o_{i}|q)}\hat{A}^{\mathrm{PO}}(q,o_{i}),$ (9)
where $G_{t}$ denotes the group size at iteration $t$ (for PPO and ReMax, $G_{t}=1$; for GRPO and RLOO, $G_{t}$ corresponds to the group size), $\pi_{\theta}$ is the current policy, $\pi_{\theta_{\operatorname{old}}}$ is the policy from the previous iteration, and $\hat{A}^{\mathrm{PO}}(q,o_{i})$ is the estimated advantage for output $o_{i}$ given query $q$.
3.3 Online Gradient Ascent
We consider online gradient ascent as our training algorithm. Specifically, at each step $t$, we randomly sample a query $q\sim D$ and generate an output $\{o_{i}\}_{i=1}^{G_{t}}\sim\pi_{\theta}(\cdot|q)$. The advantage is then computed based on the given reward function $F(q,o_{i})$, following the specific formulation of each algorithm. The gradient ascent update is performed as follows:
$\theta_{t+1}=\theta_{t}+\eta_{t}\nabla_{\theta}J_{\operatorname{PO}}(\theta_{t}).$ (10)
Assume $\pi_{\theta_{\text{old}}}=\pi_{\theta}$ for simplification, and we can write the gradient as
$\nabla_{\theta}[J_{\operatorname{PO}}(\theta)]$ $=\nabla_{\theta}\Bigg[\mathbb{E}_{q\sim D,\{o_{i}\}_{i=1}^{G_{t}}\sim\pi_{\theta_{\operatorname{old}}}(\cdot|q)}$ (11)
$\qquad\left(\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}\left[\frac{\pi_{\theta}(o_{i}|q)}{\pi_{\theta_{\operatorname{old}}}(o_{i}|q)}\hat{A}^{\mathrm{PO}}(q,o_{i})\right]\right)\Bigg]$
$=\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}\mathbb{E}_{q\sim D,o_{i}\sim\pi_{\theta}(\cdot|q)}$ (12)
$\qquad\left[\frac{\nabla_{\theta}[\pi_{\theta}(o_{i}|q)]}{\pi_{\theta}(o_{i}|q)}\hat{A}^{\mathrm{PO}}(q,o_{i})\right]$
$=\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}\mathbb{E}_{q\sim D,o_{i}\sim\pi_{\theta}(\cdot|q)}$ (13)
$\qquad\left[\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)\,\hat{A}^{\mathrm{PO}}(q,o_{i})\right].$
We can express the gradient ascent as
$\theta_{t+1}$ $=\theta_{t}+\eta_{t}\nabla_{\theta}[J_{PO}(\theta_{t})]$ (14)
$=\theta_{t}+\eta_{t}\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}\mathbb{E}_{q\sim D,o_{i}\sim\pi_{\theta}(\cdot|q)}$
$\qquad\left[\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)\,\hat{A}^{\mathrm{PO}}(q,o_{i})\right].$ (15)
In practice, the expectation cannot be directly obtained, and we rely on sampled data to approximate the gradient. Let us assume that at step $t$, we sample $N_{t}$ questions and $G_{t}$ outputs for the policy optimization algorithm. The gradient approximation is given by:
$\widehat{\nabla_{\theta}[J_{\operatorname{PO}}(\theta)]}$ $=\frac{1}{N_{t}}\sum_{j=1}^{N_{t}}\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}$
$\qquad\left[\nabla_{\theta}\log\pi_{\theta}(o_{i,j}|q_{j})\hat{A}^{\mathrm{PO}}(q_{j},o_{i,j})\right].$ (16)
The gradient ascent update rule can then be expressed as:
$\theta_{t+1}$ $=\theta_{t}+\eta_{t}\widehat{\nabla_{\theta}[J_{\operatorname{PO}}(\theta)]}$ (17)
$=\theta_{t}+\eta_{t}\frac{1}{N_{t}}\sum_{j=1}^{N_{t}}\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}$
$\qquad\left[\nabla_{\theta}\log\pi_{\theta}(o_{i,j}|q_{j})\hat{A}^{\mathrm{PO}}(q_{j},o_{i,j})\right].$ (18)
3.4 Assumptions
For the simplicity of theoretical analysis, we require the assumption as below.
Assumption 1.
Let $o\sim\pi_{\theta}(\cdot|q)$ be an output sampled from the policy for a given query $q$. We assume that the advantage is computed as
$\hat{A}^{\mathrm{PO}}(q,o)=F(q,o)-b_{\theta}(q),$ (19)
where $b_{\theta}(q)$ denotes a reference value that approximates the expected reward $\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}[F(q,o)]$. For theoretical analysis, we treat $b_{\theta}(q)$ as fixed and independent of the sampled output $o$.
Under Assumption 1, we have
$\nabla_{\theta}[J_{\operatorname{PO}}(\theta)]$
$=$ $\mathbb{E}_{q\sim D,o\sim\pi_{\theta}(\cdot|q)}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)F(q,o)\right]$
$\quad-\mathbb{E}_{q\sim D,o\sim\pi_{\theta}(\cdot|q)}\left[b_{\theta}(q)\cdot\nabla_{\theta}\log\pi_{\theta}(o|q)\right]$
$=$ $\nabla_{\theta}[J(\theta)]-\mathbb{E}_{q\sim D}\left[b_{\theta}(q)\cdot\int\nabla_{\theta}\,\pi_{\theta_{t}}(o|q)\,do\right]$
$=$ $\nabla_{\theta}[J(\theta)]-\mathbb{E}_{q\sim D}\left[b_{\theta}(q)\cdot\nabla_{\theta}\left[\int\pi_{\theta_{t}}(o|q)\,do\right]\right]$
$=$ $\nabla_{\theta}[J(\theta)]-\mathbb{E}_{q\sim D}\left[b_{\theta}(q)\cdot\nabla_{\theta}[1]\right]$
$=$ $\nabla_{\theta}[J(\theta)].$ (20)
Thus, in the following analysis, we can use $\nabla_{\theta}J(\theta)$ instead of $\nabla_{\theta}J_{\operatorname{PO}}(\theta)$. To simplify, we also use $\widehat{\nabla_{\theta}J(\theta)}$ instead of $\widehat{\nabla_{\theta}J_{\operatorname{PO}}(\theta)}$.
Table 1: Satisfaction of Assumption 1 across different algorithms.
Algorithm Assumption 1 satisfied
PPO [29]
GRPO [30]
ReMax [21]
RLOO [1]
OBLR-PO (Ours)
Assumption 2.
The logarithmic likelihood function $\log\pi_{\theta}(o|q)$ is $L$-smooth with respect to $\theta$ for all queries $q\in D$ and outputs $o$, i.e. $\forall q,o,\theta,\theta^{\prime}$,
$\|\nabla_{\theta^{\prime}}[\log\pi_{\theta^{\prime}}(o|q)]-\nabla_{\theta}[\log\pi_{\theta}(o|q)]\|_{2}\leq L\|\theta^{\prime}-\theta\|_{2}.$
Assumption 3.
We assume that there exists a uniform upper bound for the squared norm of the gradient of the log-likelihood, i.e.,
$\int\sup_{\theta}\|\nabla_{\theta}[\log\pi_{\theta}(o|q)]\|_{2}^{2}\,\mathrm{d}o\mathop{}\!\mathrm{d}q\leq M.$ (21)
Assumption 4.
The reward function is bounded, i.e., there exists a constant $B$ such that
$|F(q,o)|\leq B\quad\text{and}\quad|b_{\theta}(q)|\leq B.$ (22)
4 Main result
4.1 Bias and Variance Analysis
In this section, we prove that the gradient estimator is unbiased and has a tractable variance.
Theorem 1 (Unbiasedness).
The approximate gradient $\widehat{\nabla_{\theta}[J(\theta)]}$ is an unbiased estimator, i.e.,
$\mathbb{E}\left[\widehat{\nabla_{\theta}[J(\theta)]}\right]=\nabla_{\theta}[J(\theta)].$ (23)
We can express the approximate gradient as the sum of the true gradient and a noise term, i.e.,
$\widehat{\nabla_{\theta}[J(\theta)]}=\nabla_{\theta}[J(\theta)]+\xi(\theta),$ (24)
where $\mathbb{E}[\xi(\theta)]=0$.
Next, we analyze the covariance of the noise term $\xi(\theta)$. To simplify the expression, we define the single-sample covariance under $\pi_{\theta}(\cdot|q)$ as
$\boldsymbol{H}(\theta):=\operatorname{Var}\Big[\nabla_{\theta}\log\pi_{\theta}(o|q)\big(F(q,o)-b_{\theta}(q)\big)\Big].$ (25)
Here $o$ and $o^{\prime}$ denote two distinct samples (typically i.i.d.) from $\pi_{\theta}(\cdot|q)$. The cross-sample covariance is
$\boldsymbol{C}(\theta)=\operatorname{Cov}\Big[$ $\nabla_{\theta}\log\pi_{\theta}(o|q)\big(F(q,o)-b_{\theta}(q)\big),$
$\nabla_{\theta}\log\pi_{\theta}(o^{\prime}|q)\big(F(q,o^{\prime})-b_{\theta}(q)\big)\Big].$ (26)
Theorem 2 (Variance Expression).
The covariance matrix of $\xi(\theta)$ is given by
$\text{Var}[\xi(\theta)]=\frac{1}{N_{t}G_{t}}\boldsymbol{H}(\theta)+\frac{G_{t}-1}{N_{t}G_{t}}\boldsymbol{C}(\theta).$ (27)
The proofs of this subsection can be found in Appendix A.1.
4.2 Deriving an Upper Bound for the Loss Function
In this section, we denote the loss function as $\mathscr{L}(\theta)=J(\theta^{*})-J(\theta)$, where $\theta^{*}$ is the optimal parameter. We then derive an upper bound for $J(\theta^{*})-J(\theta_{T})$ in terms of the learning rate $\{\eta_{t}\}_{t=0}^{T-1}$, the number of queries $\{N_{t}\}_{t=0}^{T-1}$, and the group size $\{G_{t}\}_{t=0}^{T-1}$. It is clear that
$\min_{\theta}\mathscr{L}(\theta)=\mathscr{L}(\theta^{*})=0,$ (28)
indicating that the loss function attains its minimum value when $\theta=\theta^{*}$, where $L(\theta)$ vanishes.
Our main result is stated in the following theorem, whose proof can be found in Appendix A.2.
Theorem 3 (Upper Bound).
Under Assumptions 1, 2, 3, and 4, we have
$\mathbb{E}[\mathscr{L}(\theta_{T})]$ $\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}$
$\quad+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}$
$\quad+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\frac{\eta_{t}^{2}}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t})).$ (29)
4.3 Optimal Learning Rate Schedule
In this section, we aim to identify the optimal learning rate schedule $\{\eta_{t}\}_{t=0}^{T-1}$ that minimizes the final expected loss $\mathbb{E}[\mathscr{L}(\theta_{T})]$, formally stated as
$\min_{\{\eta_{t}\}_{t=0}^{T-1}}\ \mathbb{E}\left[\mathscr{L}(\theta_{T})\right].$ (30)
However, since the exact evaluation of $\mathbb{E}[\mathscr{L}(\theta_{T})]$ is generally intractable, we instead consider minimizing the upper bound provided in Theorem 3. This leads to the following result.
Theorem 4 (Optimal Learning Rate Schedule).
The optimal learning rate schedule is given by
$\eta_{t}$ $=\frac{1}{BL+B^{2}M}\cdot\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}$ (31)
$=\frac{1}{BL+B^{2}M}\cdot\frac{N_{t}\operatorname{SNR}(\theta_{t})}{1+N_{t}\operatorname{SNR}(\theta_{t})}.$ (32)
Here, we introduce the concept of the signal-to-noise ratio to measure the information content of a stochastic gradient:
$\operatorname{SNR}(\theta)$ $=\frac{\mathbb{E}\|\nabla_{\theta}[J(\theta)]\|_{2}^{2}}{\mathbb{E}[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\,\hat{A}^{\mathrm{PO}}(q,o)-\nabla_{\theta}[J(\theta)]\|_{2}^{2}]}$ (33)
$=\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta)\|_{2}^{2}}{\operatorname{tr}(\boldsymbol{H}(\theta))}.$ (34)
This theorem shows that the optimal learning rate is governed by the signal-to-noise ratio (SNR) of the gradient [40, 14].
Takeaway 1 Richer information in $\theta_{t}$ allows us to trust updates more and use a larger learning rate.
By selecting the optimal learning rate schedule, we can derive the following upper bound, as stated in the theorem below:
Theorem 5.
Under the optimal learning rate schedule 4, we have
$\quad\quad\mathbb{E}[\mathscr{L}(\theta_{T})]\leq\mathbb{E}[\mathscr{L}(\theta_{0})]$
$-\sum_{t=0}^{T-1}\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{2(BL+B^{2}M)\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))\right)}.$ (35)
Theorem 6 (Convergence Analysis).
Under the optimal learning rate schedule 4, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}=O\left(\frac{1}{\sqrt{T}}\right),$ (36)
where the big-$O$ notation hides constants and other problem-dependent parameters independent of $T$.
The proofs of this subsection can be found in Appendix A.3.
4.4 Optimal Baseline Design
In this section, we derive the optimal baseline $b_{\theta}(q)$ for the advantage as given by Theorem 3. Similarly, we optimize the upperbound of $\mathscr{L}(\theta_{T})$ instead of optimizing $\mathscr{L}(\theta_{T})$, and it suffices to minimize $\mathrm{tr}(\mathbf{H}(\theta))$, which corresponds to reducing the variance.
Theorem 7.
The optimal baseline is given by
$b_{\theta}(q)=\frac{\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\left[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}F(q,o)\right]}{\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\left[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}\right]}.$ (37)
This expression reveals that the optimal baseline depends on the gradient magnitudes rather than uniform averaging, leading to the following insight, which is consistent with observations reported in prior work [12].
Takeaway 2 The baseline should not be a simple average, but a gradient-weighted one.
The proofs of this subsection can be found in Appendix A.4.
5 Methodology
Motivated by our theoretical analysis, we propose Optimal Baseline and Learning-Rate Policy Optimization (OBLR-PO), which integrates an adaptive learning-rate schedule with an optimally designed baseline. At each iteration, the algorithm jointly adjusts the step size and reference value to reduce variance and guarantee efficient convergence.
At each iteration, we sample $G_{t}$ outputs $\{o_{i}\}_{i=1}^{G_{t}}$ from the old policy with associated rewards $F(q,o_{i})$. We first use all $G_{t}$ samples to estimate the adaptive learning rate
$\hat{\eta}_{t}=\eta_{0}\frac{N_{t}\widehat{\operatorname{SNR}(\theta_{t})}}{1+N_{t}\widehat{\operatorname{SNR}(\theta_{t})}}$, (38)
where
$\quad\quad\widehat{\operatorname{SNR}(\theta_{t})}=\frac{\|\hat{\mu}_{t}\|_{2}^{2}}{\hat{\sigma}_{t}^{2}},$ (39)
$\hat{\mu}_{t}=\frac{1}{N_{t}G_{t}}\sum_{j=1}^{N_{t}}\sum_{i=1}^{G_{t}}\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)\,F(q,o_{i}),$ (40)
$\hat{\sigma}_{t}^{2}=\frac{1}{N_{t}G_{t}-1}\sum_{j=1}^{N_{t}}\sum_{i=1}^{G_{t}}\|\nabla_{\theta}\log\pi_{\theta}(o_{i}|q)\,F(q,o_{i})-\hat{\mu}_{t}\|_{2}^{2}.$ (41)
Then, for each output $o_{i}$, we use the remaining $G_{t}-1$ samples to estimate the optimal baseline
$\hat{b}_{\theta}(q,o_{i})=\frac{\frac{1}{G_{t}-1}\sum_{j\neq i}\|\nabla_{\theta}\log\pi_{\theta}(o_{j}|q)\|_{2}^{2}F(q,o_{j})}{\frac{1}{G_{t}-1}\sum_{j\neq i}\|\nabla_{\theta}\log\pi_{\theta}(o_{j}|q)\|_{2}^{2}}$ (42)
Finally, we combine these components to construct the gradient estimator and update the policy. The complete procedure for computing the hybrid advantage and updating the policy is summarized in Algorithm 1.
Algorithm 1 OBLR-PO Update Step
0: Group rollouts $\{(q,\{o_{i},r_{i}\}_{i=1}^{G_{t}})\}$ from old policy
1: Compute the signal-to-noise ratio $\widehat{\mathrm{SNR}(\theta_{t})}$ using all $G_{t}$ samples
2: Compute the adaptive learning rate $\hat{\eta}_{t}$ using $\widehat{\mathrm{SNR}(\theta_{t})}$
3:for $i=1,\dots,G_{t}$ do
4:  For each output $o_{i}$, estimate the optimal baseline $\hat{b}_{\theta}(q,o_{i})$ using the remaining $G_{t}-1$ samples
5:  Compute the advantage $\hat{A}(q,o_{i})=r_{i}-\hat{b}_{\theta}(q,o_{i})$
6:end for
7: Construct the gradient estimator $\hat{g}_{t}$ with $\hat{\eta}_{t}$ and $\hat{A}(q,o_{i})$
8: Update policy parameters $\theta\leftarrow\theta+\hat{\eta}_{t}\cdot\hat{g}_{t}$
6 Experiments
In this section, we present a comprehensive overview of our experimental setup, evaluation metrics, and results. We evaluate the proposed Optimal Baseline and Learning-Rate Policy Optimization (OBLR-PO) algorithm on the Qwen3-4B-Base and Qwen3-8B-Base models [38], comparing its performance against the Group Relative Policy Optimization (GRPO) baseline. All experiments are implemented with the VERL framework [31] and conducted on four H200 GPUs to enable large-scale training. The subsequent subsections detail the training configuration, evaluation protocols, and empirical results.
Training Details
We conducted our experiments on two large-scale language models, Qwen3-4B-Base and Qwen3-8B-Base, employing reinforcement learning for post-training. The models were trained using our OBLR-PO algorithm as well as other RL algorithms demonstrated in Section 3.2, with the following hyperparameters:
  • Learning Rate ($\eta_{t}$): The adaptive learning rate is computed at each step based on the signal-to-noise ratio (SNR), as outlined in Theorem 4 and Algorithm 1, with an initial learning rate of $1\times 10^{-2}$.
  • Group Size ($G_{t}$): The number of outputs sampled at each step was set to $G_{t}=8$, allowing us to compare the benefits of group-based reward comparisons versus individual sampling.
  • Batch Size ($N_{t}$): The number of queries sampled at each iteration was set to $N_{t}=128$ to ensure diversity in the training set while maintaining computational efficiency.
  • Training Steps: A total of 60 training steps were conducted, with performance evaluated at each step.
Evaluations
To evaluate the effectiveness of our method, we benchmark the models on five widely used mathematical reasoning datasets:
  • OlympiadBench [13]: A benchmark of olympiad-level bilingual multimodal scientific problems, designed to evaluate high-difficulty reasoning, cross-modal understanding, and advanced problem-solving capabilities of language models.
  • GSM8K [8]: A benchmark of grade-school level math word problems, designed to measure arithmetic reasoning and step-by-step problem-solving skills of language models.
  • AIME25 [42]: A benchmark consisting of problems from the 2025 American Invitational Mathematics Examination (AIME), designed to assess advanced mathematical reasoning, precise numeric computation, and multi-step problem-solving abilities of language models.
  • MATH500 [15, 22]: A benchmark consisting of 500 high-difficulty competition-style mathematics problems from the MATH [15] dataset, designed to evaluate symbolic reasoning, multi-step deduction, and advanced mathematical problem-solving abilities of language models.
  • AMC23 [23]: A dataset derived from the 2023 American Mathematics Competition (AMC), which contains more challenging problems than GSM8K. It is commonly used to assess advanced mathematical reasoning and symbolic manipulation.
In Table 2, we report Pass@1 accuracy on all datasets, which measures the proportion of problems for which the model’s first generated answer is correct. This metric directly reflects the model’s mathematical reasoning ability and stability under reinforcement learning post-training.
Results
Table 2: Performance comparison of OBLR-PO (ours) and other baselines on five validation datasets. Bold indicates the best performance, and underline indicates the second-best.
  Model Algorithm OlympiadBench GSM8K AIME25 MATH500 AMC23
  GRPO 23.5% 88.3% 7.4% 67.6% 47.5%
PPO 22.0% 86.7% 3.7% 59.2% 47.5%
Qwen3-4B-Base ReMax 22.4% 88.0% 3.7% 65.6% 50.0%
RLOO 22.6% 87.7% 3.7% 67.8% 42.5%
OBLR-PO (Ours) 24.1% 88.6% 7.4% 67.8% 55.0%
  GRPO 26.0% 90.4% 3.7% 69.6% 42.5%
PPO 23.8% 87.0% 11.1% 64.6% 52.5%
Qwen3-8B-Base ReMax 25.3% 89.3% 7.4% 69.8% 47.5%
RLOO 24.7% 89.0% 11.1% 68.8% 55.0%
OBLR-PO (Ours) 26.0% 90.8% 11.1% 70.4% 47.5%
Figure 1: The left figure shows the Advantage during training for the Qwen3-4B-Base model, comparing GRPO (blue) and OBLR-PO (orange). The right figure shows the Advantage during training for the Qwen3-8B-Base model, again comparing GRPO (blue) and OBLR-PO (orange).
Figure 2: The left figure shows the Gradient Norm during training for the Qwen3-4B-Base model, comparing GRPO (blue) and OBLR-PO (orange). The right figure shows the Gradient Norm during training for the Qwen3-8B-Base model, again comparing GRPO (blue) and OBLR-PO (orange).
Figure 3: The left figure shows the Loss across training steps for Qwen3-4B-Base, comparing GRPO (blue) and OBLR-PO (orange). The right figure shows the Loss across training steps for Qwen3-8B-Base, again comparing GRPO (blue) and OBLR-PO (orange).
We present the results of our experiments comparing the performance of the OBLR-PO algorithm and the GRPO algorithm on the Qwen3-4B-Base and Qwen3-8B-Base models. The following metrics were analyzed during training. Additional comparisons with more reinforcement learning algorithms are provided in Appendix C.
  • Accuracy: Table 2 summarizes accuracy across five reasoning datasets after training. OBLR-PO achieves the strongest overall performance among all RL algorithms on both Qwen3-4B-Base and Qwen3-8B-Base. For Qwen3-4B-Base, OBLR-PO sets the best accuracy on every dataset, including substantial improvements such as 24.1% on OlympiadBench and 55.0% on AMC23. For Qwen3-8B-Base, OBLR-PO also achieves the best or tied-best performance on four out of five benchmarks—most notably 90.8% on GSM8K and 70.4% on MATH500. Its consistent outstanding performance across the datasets highlights its strong generalization and robustness compared to existing RL approaches.
  • Advantage: Figure 1 shows the advantage curves for the Qwen3-4B-Base and Qwen3-8B-Base models. In both cases, OBLR-PO (orange) consistently maintains a much higher and more stable advantage than GRPO (blue), leading to more favorable optimization behavior.
  • Gradient Norm: Figure 2 shows the gradient norms for Qwen3-4B-Base and Qwen3-8B-Base. OBLR-PO (orange) consistently produces lower gradients than GRPO (blue), indicating more stable training.
  • Loss: Figure 3 shows the loss across training steps for both the Qwen3-4B-Base and Qwen3-8B-Base models. OBLR-PO (orange) exhibits a much smoother and more stable loss curve compared to GRPO (blue), particularly after the initial training steps. This indicates that OBLR-PO leads to a more stable training process, with fewer fluctuations in the loss, ensuring more reliable convergence.
In summary, these results show that OBLR-PO consistently outperforms GRPO and other algorithms across multiple metrics, demonstrating enhanced stability and more favorable optimization dynamics in large-scale post-training for both the Qwen3-4B-Base and Qwen3-8B-Base models.
7 Conclusion and Limitation
In this work, we developed a theoretical framework that rigorously characterizes the bias, variance, and convergence of policy optimization under mild assumptions. Our analysis establishes the optimal learning rate schedule, governed by the signal-to-noise ratio and amplified by sample breadth and depth, and identifies the gradient-weighted baseline as a principled solution for variance reduction. These findings close the long-standing gap between heuristic algorithmic approaches and rigorous mathematical guarantees. Building on these insights, we instantiate the framework into the OBLR-PO algorithm, which consistently demonstrates stability and performance gains in large-scale post-training.
However, three limitations remain. First, our guarantees are given with respect to an upper bound on the loss, leaving a gap to the realized optimization dynamics. Second, the L-smoothness assumption (Assumption 2), while common in theory, may not strictly hold in practice and requires further empirical validation. Third, our analysis does not explicitly account for the impact of KL divergence, leaving open questions about its theoretical role in shaping optimization and generalization. We hope these findings motivate future work to tighten theoretical bounds and test assumptions in large-scale RL for LLMs.
Appendix
Appendix A Proofs for Section 4
A.1 Proofs for Section 4.1
Theorem A.1 (Unbiasedness).
The approximate gradient $\widehat{\nabla_{\theta}[J(\theta)]}$ is an unbiased estimator, i.e.,
$\mathbb{E}\left[\widehat{\nabla_{\theta}[J(\theta)]}\right]=\nabla_{\theta}[J(\theta)].$ (43)
Proof.
By the definition of the approximate gradient estimator, we have
$\widehat{\nabla_{\theta}[J(\theta)]}$ $=\frac{1}{N_{t}}\sum_{j=1}^{N_{t}}\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}\left[\nabla_{\theta}\log\pi_{\theta}(o_{i,j}|q_{j})\,\hat{A}^{\mathrm{PO}}(q_{j},o_{i,j})\right].$ (44)
Since each $q_{j}$ and its corresponding $o_{i,j}$ are sampled independently according to the policy $\pi_{\theta}$, we have
$\mathbb{E}\left[\widehat{\nabla_{\theta}[J(\theta)]}\right$] $=\mathbb{E}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\,\hat{A}^{\mathrm{PO}}(q,o)\right]=\nabla_{\theta}[J(\theta)].$ (45)
Recall that the noise term is
$\xi(\theta)=\widehat{\nabla_{\theta}J(\theta)}-\nabla_{\theta}J(\theta),$ (46)
and we use $\boldsymbol{H}(\theta)$ and $\boldsymbol{C}(\theta)$ to denote the (single-sample) variance and the cross-sample covariance, respectively:
$\boldsymbol{H}(\theta)=\operatorname{Var}\!\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\big(F(q,o)-b_{\theta}(q)\big)\right],$ (47)
$\boldsymbol{C}(\theta)=\operatorname{Cov}\!\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\big(F(q,o)-b_{\theta}(q)\big),\,\nabla_{\theta}\log\pi_{\theta}(o^{\prime}|q)\big(F(q,o^{\prime})-b_{\theta}(q)\big)\right].$ (48)
We now proceed to prove Theorem 2.
Theorem A.2 (Variance Expression).
The variance matrix of $\xi(\theta)$ is given by
$\text{Var}[\xi(\theta)]=\frac{1}{N_{t}G_{t}}\boldsymbol{H}(\theta)+\frac{G_{t}-1}{N_{t}G_{t}}\boldsymbol{C}(\theta).$ (49)
Proof.
By the definition of $\xi(\theta)$, we have
$\text{Var}[\xi(\theta)]$ $=\text{Var}\left[\frac{1}{N_{t}}\sum_{j=1}^{N_{t}}\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}\left[\nabla_{\theta}\log\pi_{\theta}(o_{i,j}|q_{j})\hat{A}^{\mathrm{PO}}(q_{j},o_{i,j})\right]-\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(o|q)\hat{A}^{\mathrm{PO}}(q,o)]\right]$ (50)
$=\frac{1}{N_{t}^{2}G_{t}^{2}}\sum_{j=1}^{N_{t}}\text{Var}\left[\sum_{i=1}^{G_{t}}\nabla_{\theta}\log\pi_{\theta}(o_{i,j}|q_{j})\hat{A}^{\mathrm{PO}}(q_{j},o_{i,j})\right]$ (51)
$=\frac{1}{N_{t}G_{t}}\text{Var}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\hat{A}^{\mathrm{PO}}(q,o)\right]$ (52)
$+\frac{G_{t}-1}{N_{t}G_{t}}\text{Cov}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\hat{A}^{\mathrm{PO}}(q,o),\nabla_{\theta}\log\pi_{\theta}(o^{\prime}|q)\hat{A}^{\mathrm{PO}}(q,o^{\prime})\right]$ (53)
$=\frac{1}{N_{t}G_{t}}\text{Var}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)(F(q,o)-b_{\theta}(q))\right]$ (54)
$+\frac{G_{t}-1}{N_{t}G_{t}}\text{Cov}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)(F(q,o)-b_{\theta}(q)),\nabla_{\theta}\log\pi_{\theta}(o^{\prime}|q)(F(q,o^{\prime})-b_{\theta}(q))\right]$ (55)
$=\frac{1}{N_{t}G_{t}}\boldsymbol{H}(\theta)+\frac{G_{t}-1}{N_{t}G_{t}}\boldsymbol{C}(\theta).$ (56)
A.2 Proofs for Section 4.2
Theorem A.3 (Upper Bound).
Under Assumptions 1, 2, 3, and 4, we have
$\mathbb{E}[\mathscr{L}(\theta_{T})]\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$ (57)
Lemma A.4.
Let $f(x,\theta)\colon\mathscr{X}\times\Theta\to\mathbb{R}$ be a nonnegative binary function that is integrable with respect to $x$ and differentiable with respect to $\theta$. Assume the following conditions:
$\int_{\mathscr{X}}f(x,\theta)\,\mathrm{d}x=1,$ (58)
for all $x\in\mathscr{X}$,
$\|\nabla_{\theta}[\log f(x,\theta^{\prime})]-\nabla_{\theta}[\log f(x,\theta)]\|_{2}\leq L\|\theta^{\prime}-\theta\|_{2},$ (59)
and that there exists a uniform upper bound for $\int_{\mathscr{X}}\nabla_{\theta}[\log f(x,\theta)]\,\mathrm{d}x$, i.e.,
$\int_{\mathscr{X}}\sup_{\theta\in\Theta}\|\nabla_{\theta}[\log f(x,\theta)]\|_{2}^{2}\,\mathrm{d}x\leq M.$ (60)
Then, it follows that
$\int_{\mathscr{X}}\|\nabla_{\theta}f(x,\theta^{\prime})-\nabla_{\theta}f(x,\theta)\|_{2}\,\mathrm{d}x\leq(L+M)\|\theta^{\prime}-\theta\|_{2}.$ (61)
Proof of Lemma A.4.
By the chain rule, we have
$\nabla_{\theta}[f(x,\theta)]=\nabla_{\theta}[\log f(x,\theta)]\cdot f(x,\theta).$ (62)
Then we substitute it into (62) and apply the triangle inequality,
$\int_{\mathscr{X}}\|\nabla_{\theta}f(x,\theta^{\prime})-\nabla_{\theta}f(x,\theta)\|_{2}\,\mathrm{d}x$ (63)
$\leq$ $\int_{\mathscr{X}}\|\nabla_{\theta}[\log f(x,\theta^{\prime})]\cdot f(x,\theta^{\prime})-\nabla_{\theta}[\log f(x,\theta)]\cdot f(x,\theta)\|_{2}\,\mathrm{d}x$ (64)
$\leq$ $\underbrace{\int_{\mathscr{X}}f(x,\theta)\|\nabla_{\theta}\log[f(x,\theta^{\prime})]-\nabla_{\theta}\log[f(x,\theta)]\|_{2}\mathop{}\!\mathrm{d}x}_{\operatorname{I}}+\underbrace{\int_{\mathscr{X}}|f(x,\theta^{\prime})-f(x,\theta)|\|\nabla_{\theta}\log[f(x,\theta^{\prime})]\|_{2}\mathop{}\!\mathrm{d}x}_{\operatorname{II}}.$ (65)
Then we bound $\operatorname{I}$ and $\operatorname{II}$ separately.
To bound the first term, we just need to apply the conditions and have
$\operatorname{I}\leq L\|\theta^{\prime}-\theta\|_{2}\int_{\mathscr{X}}f(x,\theta)\mathop{}\!\mathrm{d}x=L\|\theta^{\prime}-\theta\|_{2}.$ (66)
For the second term, we have
$\operatorname{II}$ $\leq\|\theta^{\prime}-\theta\|_{2}\int_{\mathscr{X}}\|\nabla_{\theta}[f(x,\tilde{\theta}(x))]\|\cdot\|\nabla_{\theta}\log[f(x,\theta^{\prime})]\|_{2}\mathop{}\!\mathrm{d}x$ (67)
$\leq\|\theta^{\prime}-\theta\|_{2}\int_{\mathscr{X}}B(x)^{2}\mathop{}\!\mathrm{d}x$ (68)
$\leq M\|\theta^{\prime}-\theta\|_{2},$ (69)
where $B(x):=\sup_{\theta}f(x,\theta)$.
Combine them together, and we prove the Lemma A.4. ∎
Then we come back to Theorem A.3.
Proof of Theorem A.3.
Apply the Lemma A.4 to $F(q,o)\pi_{\theta}(o|q)$, we have
$\|\nabla_{\theta}[F(q,o)\log\pi_{\theta^{\prime}}(o|q)]-\nabla_{\theta}[F(q,o)\log\pi_{\theta}(o|q)]\|$ (70)
$\leq$ $|F(q,o)|\|\nabla_{\theta}[\log\pi_{\theta^{\prime}}(o|q)]-\nabla_{\theta}[\log\pi_{\theta}(o|q)]\|$ (71)
$\leq$ $BL\|\theta^{\prime}-\theta\|_{2},$ (72)
and
$\int\sup_{\theta\in\Theta}\|\nabla_{\theta}[F(q,o)\log\pi_{\theta}(o|q)]\|_{2}^{2}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}q\leq F(q,o)^{2}\int\sup_{\theta\in\Theta}\|\nabla_{\theta}[\log\pi_{\theta}(o|q)]\|_{2}^{2}\mathop{}\!\mathrm{d}o\mathop{}\!\mathrm{d}q\leq B^{2}M.$ (73)
Thus, we have for fixed $q$,
$\int\|\nabla_{\theta}[\pi_{\theta^{\prime}}(o|q)F(q,o)]-\nabla_{\theta}[\pi_{\theta}(o|q)F(q,o)]\|\mathop{}\!\mathrm{d}o\leq(BL+B^{2}M)\|\theta^{\prime}-\theta\|_{2}.$ (74)
So we can identify the smoothness coefficient of $\mathscr{L}(\theta)$ as follows.
$\|\nabla_{\theta}J(\theta^{\prime})-\nabla_{\theta}J(\theta)\|_{2}$ $\leq\|\mathbb{E}_{q\sim D}[\nabla_{\theta}[\mathbb{E}_{o\sim\pi_{\theta^{\prime}}(\cdot|q)}F(q,o)]-\nabla_{\theta}[\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}F(q,o)]]\|_{2}$ (75)
$\leq\mathbb{E}_{q\in D}\left[\int\|\nabla_{\theta}[\pi_{\theta^{\prime}}(o|q)F(q,o)]-\nabla_{\theta}[\pi_{\theta}(o|q)F(q,o)]\|\mathop{}\!\mathrm{d}o\right]$ (76)
$\leq(BL+B^{2}M)\|\theta^{\prime}-\theta\|_{2},$ (77)
which implies that $J(\theta)$ is $(BL+B^{2}M)$-smooth.
Then apply the Taylor’s expansion,
$\mathscr{L}(\theta_{t+1})\leq\mathscr{L}(\theta_{t})-\langle\nabla_{\theta}\mathscr{L}(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{(BL+B^{2}M)}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}.$ (78)
Note that $\xi(\theta_{t})$ is independent of $\theta_{t}$, and thus we have
$\mathbb{E}[\mathscr{L}(\theta_{t+1})]\leq\mathbb{E}[\mathscr{L}(\theta_{t})]-\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{\eta_{t}^{2}(BL+B^{2}M)}{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\operatorname{tr}(\operatorname{Var}(\xi(\theta_{t})))\right).$ (79)
It remains to control $\operatorname{tr}(\operatorname{Var}(\xi(\theta_{t})))$. Recall that $\xi(\theta_{t})$ is an average of $N_{t}$ independent queries, each containing $G_{t}$ samples $o_{1},\dots,o_{G_{t}}$ drawn under the same $q$, so the variance decomposes into a within-sample term and a cross-sample term:
$\operatorname{tr}(\operatorname{Var}(\xi(\theta_{t})))=\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))+\frac{G_{t}-1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{C}(\theta_{t})).$ (80)
Let $X:=\nabla_{\theta}\log\pi_{\theta}(o|q)\big(F(q,o)-b_{\theta}(q)\big)$ and $Y:=\nabla_{\theta}\log\pi_{\theta}(o^{\prime}|q)\big(F(q,o^{\prime})-b_{\theta}(q)\big)$, where $o$ and $o^{\prime}$ are distinct samples under the same $q$. For any unit vector $u$, Cauchy–Schwarz gives
$u^{\top}\operatorname{Cov}(X,Y)u=\operatorname{Cov}(u^{\top}X,\,u^{\top}Y)\leq\sqrt{\operatorname{Var}(u^{\top}X)\operatorname{Var}(u^{\top}Y)}=\operatorname{Var}(u^{\top}X),$ (81)
where the last equality uses that $X$ and $Y$ are identically distributed (given $q$). Hence $\operatorname{Cov}(X,Y)\preceq\operatorname{Var}(X)$, implying $\operatorname{tr}(\boldsymbol{C}(\theta_{t}))\leq\operatorname{tr}(\boldsymbol{H}(\theta_{t}))$. Therefore,
$\operatorname{tr}(\operatorname{Var}(\xi(\theta_{t})))\leq\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))+\frac{G_{t}-1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))=\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t})).$ (82)
Thus, we have
$\mathbb{E}[\mathscr{L}(\theta_{t+1})]\leq\mathbb{E}[\mathscr{L}(\theta_{t})]-\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{\eta_{t}^{2}(BL+B^{2}M)}{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$ (83)
Summing over $t=0,\dots,T-1$, we obtain
$\mathbb{E}[\mathscr{L}(\theta_{T})]\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$ (84)
A.3 Proofs for Section 4.3
Theorem A.5 (Optimal Learning Rate Schedule).
The optimal learning rate schedule is given by
$\eta_{t}=\frac{1}{BL+B^{2}M}\cdot\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}=\frac{1}{BL+B^{2}M}\cdot\frac{N_{t}\operatorname{SNR}(\theta_{t})}{1+N_{t}\operatorname{SNR}(\theta_{t})}.$ (85)
Here, we introduce the concept of the signal-to-noise ratio to measure the information content of a stochastic gradient:
$\operatorname{SNR}(\theta)=\frac{\mathbb{E}\|\nabla_{\theta}[J(\theta)]\|_{2}^{2}}{\mathbb{E}[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\,\hat{A}^{\mathrm{PO}}(q,o)-\nabla_{\theta}[J(\theta)]\|_{2}^{2}]}=\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta)\|_{2}^{2}}{\operatorname{tr}(\boldsymbol{H}(\theta))}.$ (86)
Proof.
From the upper bound in Equation (57), each term involving $\eta_{t}$ takes the form of a quadratic function:
$\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\eta_{t}^{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$
This is a convex quadratic function in $\eta_{t}$, and its minimum is achieved by setting the derivative to zero. Solving for the optimal $\eta_{t}$ gives the expression in Theorem 4. ∎
Theorem A.6.
Under the optimal learning rate schedule in Theorem 4, we have
$\mathbb{E}[\mathscr{L}(\theta_{T})]\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{2(BL+B^{2}M)\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))\right)}.$ (87)
Proof.
Substitute the optimal learning rate in Theorem 4, we have
$\mathbb{E}[\mathscr{L}(\theta_{T})]$ $\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)$ (88)
$=\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}\cdot\frac{1}{2(BL+B^{2}M)}.$ (89)
Rearranging terms completes the proof. ∎
Theorem A.7 (Convergence Analysis).
Under the optimal learning rate schedule in Theorem 4, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}=O\left(\frac{1}{\sqrt{T}}\right),$ (90)
where the big-$O$ notation hides constants and other problem-dependent parameters independent of $T$.
Lemma A.8.
The trace of the covariance matrix $\boldsymbol{H}(\theta)$ is bounded.
$\operatorname{tr}(\boldsymbol{H}(\theta))\leq 4B^{2}M.$ (91)
Proof of Lemma A.8.
We have
$\operatorname{tr}(\boldsymbol{H}(\theta))$ $=\mathbb{E}[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\,(F(q,o)-b_{\theta}(q))-\nabla_{\theta}[J(\theta)]\|_{2}^{2}]$ (92)
$\leq\mathbb{E}[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\,(F(q,o)-b_{\theta}(q))\|_{2}^{2}]$ (93)
$\leq 4B^{2}\cdot\mathbb{E}[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}]$ (94)
$\leq 4B^{2}M,$ (95)
where the second inequality follows from Assumption 4, which bounds $|F(q,o)-b_{\theta}(q)|\leq 2B$, and the last inequality follows from Assumption 3, which ensures $\mathbb{E}[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}]\leq M$. ∎
Then we come back to Theorem A.7. For notational simplicity, we write $\lesssim$ (resp. $\gtrsim$) to indicate an upper (resp. lower) bound up to a constant factor independent of $T$.
Proof of Theorem A.7.
By the results of Theorem 5,
$\sum_{t=0}^{T-1}\frac{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}$ $\leq 2(BL+B^{2}M)\cdot\left(\mathbb{E}[\mathscr{L}(\theta_{0})]-\mathbb{E}[\mathscr{L}(\theta_{T})]\right)$ (96)
$\leq 2(BL+B^{2}M)\cdot\mathbb{E}[\mathscr{L}(\theta_{0})]$ (97)
By applying the Cauchy-Schwarz inequality, we have
$\left(\sum_{t=0}^{T-1}\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}\right)$ $\cdot\left(\sum_{t=0}^{T-1}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)\right)$ (98)
$\geq\left(\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}$ (99)
Combine them together and we have
$\frac{\left(\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}}{\sum_{t=0}^{T-1}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)}=O(1).$ (100)
By the Lemma A.8, we have
$1\gtrsim\frac{\left(\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}}{\sum_{t=0}^{T-1}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)}\geq\frac{\left(\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}}{\sum_{t=0}^{T-1}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)+4TB^{2}M}.$ (101)
And thus, we have
$\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\lesssim\frac{1+\sqrt{1+16TB^{2}M}}{2}\leq 1+\sqrt{4TB^{2}M}\lesssim\sqrt{T}.$ (102)
Thus, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}=O\left(\frac{1}{\sqrt{T}}\right)$ (103)
A.4 Proofs for Section 4.4
Theorem A.9.
The optimal baseline is given by
$b_{\theta}(q)=\frac{\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\left[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}F(q,o)\right]}{\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\left[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}\right]}.$ (104)
Proof.
As $\operatorname{tr}(\boldsymbol{H}(\theta))$ can be expressed as
$\operatorname{tr}(\boldsymbol{H}(\theta))$ $=\|\nabla_{\theta}\log\pi_{\theta}(o|q)(F(q,o)-b_{\theta}(q))\|_{2}^{2}$ (105)
$=\mathbb{E}_{q\sim D}\Bigg[\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\big[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}\big]b_{\theta}(q)^{2}$ (106)
$\qquad-2\,\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\big[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}F(q,o)\big]b_{\theta}(q)$ (107)
$\qquad+\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\big[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}F(q,o)^{2}\big]\Bigg].$ (108)
As $b_{\theta}(q)$ takes the form of a quadratic function, the optimal $b_{\theta}(q)$ is given by
$b_{\theta}(q)=\frac{\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\left[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}F(q,o)\right]}{\mathbb{E}_{o\sim\pi_{\theta}(\cdot|q)}\left[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\|_{2}^{2}\right]}.$ (109)
Appendix B Further Analysis
B.1 Optimal Query Sampling Strategy Under a Data Constraint
In this section, we study the optimal query sampling schedule $\{N_{t}\}_{t=0}^{T-1}$ under a fixed data budget. Formally, we consider
$\min_{\{N_{t}\}_{t=0}^{T-1},\,\{G_{t}\}_{t=0}^{T-1}}\;\mathbb{E}\big[\mathscr{L}(\theta_{T})\big]\quad\text{s.t.}\quad\sum_{t=0}^{T-1}N_{t}\leq C.$ (110)
Theorem B.1 (Optimal Sampling Strategy).
Under the computational budget constraint, the optimal sampling strategy is given by
$N_{t}=\frac{C+\sum_{t=0}^{T-1}\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}{\sum_{t=0}^{T-1}\sqrt{\text{tr}(\boldsymbol{H}(\theta_{t}))}}\sqrt{\text{tr}(\boldsymbol{H}(\theta_{t}))}-\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}.$ (111)
Proof.
To minimize $\mathbb{E}[\mathscr{L}(\theta_{T})]$, we aim to maximize the following expression:
$\sum_{t=0}^{T-1}\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))},$ (112)
which is equivalent to minimizing
$\sum_{t=0}^{T-1}\frac{\frac{1}{N_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))}=\sum_{t=0}^{T-1}\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{N_{t}+\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}.$ (113)
By applying the Cauchy-Schwarz inequality, we have
$\left(\sum_{t=0}^{T-1}\left(N_{t}+\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}\right)\right)\cdot\left(\sum_{t=0}^{T-1}\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{N_{t}+\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}\right)\geq\left(\sum_{t=0}^{T-1}\sqrt{\text{tr}(\boldsymbol{H}(\theta_{t}))}\right)^{2}.$ (114)
Equality holds when
$\frac{\sqrt{\text{tr}(\boldsymbol{H}(\theta_{t}))}}{N_{t}+\frac{\text{tr}(\boldsymbol{H}(\theta_{t}))}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}}=\operatorname{Const}.$ (115)
Substituting this into the computational budget constraint yields the result. ∎
Appendix C Additional Results
In this section, we provide additional results, including the advantages and gradient norms of other algorithms, as well as the behaviors of KL loss and entropy.
Figure 4: Advantages observed for Qwen3-4B-Base (left) and Qwen3-8B-Base (right) during training.
Figure 5: Gradient norm curves of Qwen3-4B-Base (left) and Qwen3-8B-Base (right) during training.
Figure 6: KL loss trajectories of Qwen3-4B-Base (left) and Qwen3-8B-Base (right) during training.
Figure 7: Entropy trajectories of Qwen3-4B-Base (left) and Qwen3-8B-Base (right) during training.